OpenStudy (doodleduck8):

Solve the equation for y: 4x - 2y = 6 1. y = -2x + 3 2. y = 2x + 3 3. y = -2x - 3 4. y = 2x - 3 I got 4: y = 2x - 3 is this correct, if it is tell me how you figured it out step be step

jimthompson5910 (jim_thompson5910):

Yes it's correct. 4x - 2y = 6 4x - 2y+2y = 6+2y ... add 2y to both sides 4x = 6+2y 4x-6 = 6+2y-6 ... subtract 6 from both sides 4x-6 = 2y 2y = 4x-6 y = 2x-3 ... divide every term by 2

OpenStudy (doodleduck8):

Thanks @jim_thompson5910

OpenStudy (doodleduck8):

That really helped

jimthompson5910 (jim_thompson5910):

I'm glad it helped